OpenStudy (anonymous):

How can I correct the right side? (√y +1)^2=y+1

OpenStudy (abb0t):

take the square root of everything

OpenStudy (anonymous):

Can you explain? I was foiling the left side.

OpenStudy (abb0t):

do u want the right side to equal the left side? or r u trying to solve for "x"?

OpenStudy (anonymous):

I need to fix the right side

OpenStudy (abb0t):

what do u mean by fix?

OpenStudy (anonymous):

I have to make the right side equal the left. Like right now the right side is wrong

OpenStudy (abb0t):

\[(\sqrt{x}+1)^2=y+1\] \[\sqrt{(\sqrt{x}+1)^2}= (\sqrt{x}+1)\] does that make it easier?

OpenStudy (anonymous):

I don't understand how you got x's

OpenStudy (anonymous):

first take the square root both side then square them again you will have y+1=y+1